P7-3 Bond Prices [LO2]

Staind, Inc., has 7 percent coupon bonds on the market that have 10 years left to maturity. The bonds make annual payments. If the YTM on these bonds is 9 percent, what is the current bond price?


rev: 09_18_2012

Multiple Choice

  • $915.23

  • $828.06

  • $1,144.47

  • $1,030.00

  • $871.65
Answer #1

Annual coupon=1000*7%=70

Hence current price=Annual coupon*Present value of annuity factor(9%,10)+$1000*Present value of discounting factor(9%,10)
=70*6.417657701+$1000*0.422410806

=$871.65(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=70[1-(1.09)^-10]/0.09

=70*6.417657701

2.Present value of discounting factor=1000/1.09^10

=1000*0.422410806